OpenStudy (anonymous):

Can someone tell me if I got these right or wrong please. 1.Choose the point-slope form of the equation below that represents the line that passes through the point (−1, 6) and has a slope of −3. Is the answer y-6=3(x+1) 2.Choose the point-slope form of the equation below that represents the line that passes through the points (−6, 4) and (2, 0). Is the answer y-4= negative one half(x+6)

OpenStudy (anonymous):

the frist on is y + 3 =1/2 (x − 6)

OpenStudy (anonymous):

So is the second one right

OpenStudy (anonymous):

hold on kk

OpenStudy (anonymous):

First Of All You Need To Find The Slope (x1,y1)=(-6, 4) And (x2,y2)= (2, 0).

OpenStudy (anonymous):

y - 0= -1/2 (x - 2)